A flint glass prism and a crown glass prism are to be combined in such a way that the deviation of the mean ray is zero. The refractive index of flint and crown glasses for the mean ray are 1.620 and 1.518 respectively. If the refracting angle of the flint prism is 6.0°, what would be the refracting angle of the crown prism?

Grade:upto college level

2 Answers

Navjyot Kalra
9 years ago
1. Given that, Refractive index of flint glass = μ base f = 1.620 Refractive index of crown glass = μ basec = 1.518 Refracting angle of flint prism = A base f = 6.0° For zero net deviation of mean ray (μ base f – 1)Af = (μ base c – 1) Ac ⇒ A base c = μ base 1 – 1/μ base c – 1A base f = 1.620 – 1/1.518 – 1(6.0)° = 7.2°
Deepak Patra
9 years ago
Sol. Given that, Refractive index of flint glass = μ base f = 1.620 Refractive index of crown glass = μ basec = 1.518 Refracting angle of flint prism = A base f = 6.0° For zero net deviation of mean ray (μ base f – 1)Af = (μ base c – 1) Ac ⇒ A base c = μ base 1 – 1/μ base c – 1A base f = 1.620 – 1/1.518 – 1(6.0)° = 7.2°